Annie drives from town A to town B in 3 hours(at a fixed speed). The traffic is heavier on the return trip, so Annie's speed is 12 mph less and the return trip takes 4 hours. What were her speeds on both parts of the trip?
Answer by lwsshak3(11628)

Annie drives from town A to town B in 3 hours(at a fixed speed). The traffic is heavier on the return trip, so Annie's speed is 12 mph less and the return trip takes 4 hours. What were her speeds on both parts of the trip?
**
distance=speed*travel time
let x=speed of forward trip
x-3= speed of return trip
..
3x=4(x-3)
3x=4x-12
x=12
x-3=9
ans:
speed of forward trip=12 mph
speed of return trip =9 mph